OpenStudy (anonymous):

If (a^m)*(a^3)=(1/a^2), what is a? Could someone give me an explanation as well?

OpenStudy (skullpatrol):

Are you sure you are asked for a, not m?

OpenStudy (anonymous):

Unfortunately, they asked for a. If it was m, then I already know what it would be

OpenStudy (skullpatrol):

Then a could be any nonzero number.

OpenStudy (anonymous):

I see. Is there a rule that gives that away? If it was asking for m, then the exponent would be -5 right? I think there was a typo on my homework since it doesn't make much sense to ask for a.

OpenStudy (skullpatrol):

Yes, the exponent is -5.

OpenStudy (skullpatrol):

The way you know it is -5 is by using the rule of exponents which is only true if a is not zero.

OpenStudy (skullpatrol):

Simply because division by 0 is undefined.

OpenStudy (anonymous):

Alright I'm getting it now. Then a can be any number such as 2,3,4 and still be 1/a^2 then?

OpenStudy (skullpatrol):

Correct.

OpenStudy (anonymous):

Okay, thank v much for the explanation

OpenStudy (skullpatrol):

Would you like an explanation why division by 0 is undefined?

OpenStudy (anonymous):

No one really elaborated on it to me for. So yes, thank you, I would like to know actually.

OpenStudy (skullpatrol):

what grade level?

OpenStudy (anonymous):

Pre-calculus now

OpenStudy (skullpatrol):

let me explain it using the exponent rule: a^m = a*a*a*...*a right?

OpenStudy (anonymous):

Yes, that is correct.

OpenStudy (skullpatrol):

That is for m = any positive integer: 1, 2, 3, 4, ... right?

OpenStudy (skullpatrol):

What about for the negative integers: -1, -2, -3, -4,... what is the rule?

OpenStudy (skullpatrol):

$$\Huge a^{-m}=?$$

OpenStudy (anonymous):

Then the produce would have to become a fraction when a negative integer is the exponent, correct?

OpenStudy (anonymous):

product*

OpenStudy (skullpatrol):

$$\Huge a^{-m}=\dfrac{1}{a^m}$$

OpenStudy (anonymous):

Yes, just like that.

OpenStudy (skullpatrol):

Now, we have a rule for positive integers and negative integers, what about a 0 exponent?

OpenStudy (anonymous):

I'm curious though, if a or m was equal to 1, then wouldn't that change the product too?

OpenStudy (anonymous):

If 0 is the exponent then the problem would equal to 1

OpenStudy (skullpatrol):

If a or m was 1 you substitute 1 into the above formula and it will work out fine :-)

OpenStudy (anonymous):

I understand much better already. Thanks once again^^

OpenStudy (skullpatrol):

The real question becomes WHY is $$\Huge a^{0}=?=1$$for any nonzero a

OpenStudy (anonymous):

Is it not because the base is being divided by itself I wonder? If it were to have the exponent of 1 then the base would stay the same.

OpenStudy (skullpatrol):

What you are wondering is EXACTLY correct :D

OpenStudy (anonymous):

Jeez, I think I actually do not really understand how any number can become 1 just from the exponent of 0 though since it's more based on roots and and being multiplied repeatedly and no division. Is there a clearer way to explain this?

OpenStudy (anonymous):

Unless division is really behind it

OpenStudy (skullpatrol):

$$\Huge a^{0}=\dfrac{a}{a}$$

OpenStudy (anonymous):

Yep never mind, I see very clearly now

OpenStudy (skullpatrol):

Now, we can try to understand why this is true for EVERY real number except 0...ready?

OpenStudy (anonymous):

Sure, visuals help very much^^

OpenStudy (skullpatrol):

"never mind"? Ok, you explain why it doesn't work for 0 to me :-)

OpenStudy (anonymous):

0 wouldn't work in place of a(the base) since 0 to any power at all would remain zero either way. Then it would be just 0 divided by 0

OpenStudy (anonymous):

Which is undefined?

OpenStudy (anonymous):

Am I on the wrong track?

OpenStudy (skullpatrol):

Yes, you are the right track: $$\Huge 0^{0}=\dfrac{0}{0}$$

OpenStudy (skullpatrol):

for ALL the other numbers that^ equals 1, right?

OpenStudy (anonymous):

Yes.

OpenStudy (skullpatrol):

But, it equals 1 because $$\Huge\dfrac{a}{a}= a\cdot\cfrac{1}{a}$$

OpenStudy (skullpatrol):

$$\Huge\dfrac{a}{a}= a\cdot\cfrac{1}{a}=1$$

OpenStudy (anonymous):

Yes, since the a's cancel out in the center to equal 1

OpenStudy (anonymous):

Which zero could never achieve if it was a

OpenStudy (skullpatrol):

Try not to think in terms of "canceling" yet. Just take any number and multiply it by the the result of 1 divided by the same number. Like 5/5 = 5 * (1/5) = 5 * (0.2) = 1

OpenStudy (anonymous):

Oh I see. I'm sorry. Then in the end we are dividing the same number together to equal And with your example of 5 * (0.2)=1 then certain decimals work perfectly fine to make 1, but it won't come up in exponential problems will it?

OpenStudy (skullpatrol):

Yes, it does come up in the exponential problem because $$\Huge\dfrac{a}{a}= a\cdot\cfrac{1}{a}=1$$ $$\Huge\dfrac{0}{0}= 0\cdot\cfrac{1}{0}=0 $$ because 0 times any real number is 0

OpenStudy (anonymous):

Ah, I get it. (0.2) is also (1/5) which is usable for a power.

OpenStudy (skullpatrol):

Yes, but 1/0 is not "usable"

OpenStudy (anonymous):

Alright, any integer is doable besides 0 since most circumstances that regular integers go through wouldn't work, such as being divided by other numbers. I'm sorry, it must be a pain for you to write out these equations. Am I miss anything else important?

OpenStudy (skullpatrol):

Dividing by 0 would mean multiplying by 1/0 but 0 has no reciprocal because 0 times any number is 0, not 1.

OpenStudy (skullpatrol):

Therefore, division by 0 has no meaning in the set of real numbers.

OpenStudy (anonymous):

Okay, it becomes clearer now that I also keep in mind that though 0 is real number and an integer it doesn't have a reciproval. I think what you showed me was the inverse property of multiplication right?

OpenStudy (anonymous):

Oh dear, I kept you here for an hour. :S

OpenStudy (skullpatrol):

Good job :D

OpenStudy (anonymous):

Thank you, I really appreciate you writing these examples and taking your time out to help me :)

OpenStudy (skullpatrol):

$$\Huge a^0=\dfrac{a}{a}= a\cdot\cfrac{1}{a}=1$$ $$\Huge 0^0=\dfrac{0}{0}= 0\cdot\cfrac{1}{0}=0$$

OpenStudy (skullpatrol):

See^ the contradiction?

OpenStudy (anonymous):

Yes, 0 doesn't follow the same rules

OpenStudy (skullpatrol):

Because it has its OWN rule that says 0*(any number) = 0

OpenStudy (skullpatrol):

So, 1/0 cannot be (any number)

OpenStudy (anonymous):

Then a good summary would be that 0^0 or to any power would mostly be meaningless?

OpenStudy (anonymous):

Since 0/0=0 and 0*0=0 always

OpenStudy (anonymous):

Well 0/0 is considered undefined, but with 1/0 does it =0 or is undefined?

OpenStudy (skullpatrol):

The expression 0^0 has no meaning. Because division by 0 is undefined.

OpenStudy (skullpatrol):

1/0 is undefined too.

OpenStudy (anonymous):

Okay, I think I fully understand by now.